# Chain-of-Query: Unleashing the Power of LLMs in SQL-Aided Table Understanding via Multi-Agent Collaboration

### v1

@@ 3.5 Balancing SQL and LLM Reasoning through Dynamic Planning
We build the Dynamic Planner that determines whether to continue SQL generation, guided by a Sufficiency-based Early Stopping Mechanism. Together with the Answer Generator, it implements our Hybrid Reasoning Division Strategy: SQL handles mechanical reasoning, while LLMs perform logical reasoning over intermediate results to produce the final answer.
@@ 3.5.1 Hybrid Reasoning Division Strategy
Existing SQL-aided approaches treat the SQL output as the final answer, requiring highly precise queries. This often leads to long, fragile SQL programs, especially when reasoning is complex or question intent is ambiguous. However, table understanding is inherently open-ended: the goal is to derive a natural language answer, not merely to retrieve an exact value. Based on these insights, we introduce a Hybrid Reasoning Division Strategy that decomposes table understanding into two stages: mechanical reasoning (e.g., filtering, arithmetic) is offloaded to SQL, while logical reasoning (e.g., comparison, inference) is handled by LLMs. Crucially, SQL execution results are treated as intermediate data rather than final answers, enabling LLMs to apply their stronger inference capabilities over SQL outputs to arrive at the final answer. This division leverages the respective strengths of SQL and LLMs. Our case study (Appendix A) further explains this design.
3.5.2 Sufficiency-based Early Stopping
To support this hybrid reasoning strategy, we introduce the Sufficiency-based Early Stopping Mechanism. It enables the Planner to halt clause generation once the retrieved data are sufficient. This prevents unnecessary SQL complexity and ensures that reasoning responsibilities are dynamically and appropriately balanced between SQL and LLMs.
3.6 Answer Generator
Table 1: Results on WikiTQ and TabFact using GPT-3.5 and LLaMA 2. Acc. = accuracy (%), Inv. = invalid SQL rate (%, lower is better). Underline = second-best, bold = best. Improvements are over second-best. CoQ achieves the highest accuracy with a large margin, driven by schema abstraction, clause-by-clause SQL generation, and hybrid reasoning.
@@ 3.6 Answer Generator
Generic Table Understanding
@@ 4 Experiments
In this section, we empirically evaluate the effectiveness of CoQ and aim to address the following questions: RQ1: How does CoQ compare with generic table understanding methods? RQ2: How does CoQ perform relative to traditional SQL-aided methods? RQ3: How well does CoQ generalize to real-world, structurally complex tabular workloads?
@@ 4.1 Experimental Setup
Datasets and Evaluation Metrics.  We evaluate our CoQ framework on five widely used table understanding benchmarks: WikiTQ (Pasupat and Liang, 2015), FeTaQA (Nan et al., 2022), TabFact (Chen et al., 2020), IM-TQA (Zheng et al., 2023), and Open-WikiTable (Kweon et al., 2023). WikiTQ and FeTaQA are table QA datasets requiring short-span and free-form answers, respectively. TabFact is a fact verification task based on tables. IM-TQA features complex real-world table styles, while Open-WikiTable involves multi-table scenarios. Collectively, these datasets cover diverse reasoning types, table structures, and domains, forming a robust testbed for evaluating generalization. Dataset details are provided in Appendix F.
We use official accuracy metrics for WikiTQ, IM-TQA, and Open-WikiTable, and standard binary accuracy for TabFact. For FeTaQA, we report BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) scores to align with prior work. We also track the rate of invalid SQL generation.
Baselines.  We compare our method against two categories of baseline approaches: (a) Generic table understanding, including End-to-End QA, Few-Shot QA, Table-to-Text (Min et al., 2024), Chain-of-Thought (Wei et al., 2022), Binder (Cheng et al., 2023), Dater (Ye et al., 2023), Chain-of-Table (Wang et al., 2024c), and Tree-of-Table (Ji et al., 2024); (b) SQL-aided table understanding, including Basic Text-to-SQL (Rajkumar et al., 2022), OpenTab (Kong et al., 2024a), MAC-SQL (Wang et al., 2025b), and MAG-SQL (Xie et al., 2024b).
Implementation Details.  To align with baselines, we adopt GPT-3.5 and LLaMA 2 as the backbone LLMs, and additionally evaluate with the recent stronger models LLaMA 3.1, DeepSeek-V3, and GPT-4.1. Model configurations are detailed in Appendix H. Prompts include few-shot examples sampled from the training set, with illustrative cases shown in Appendix I.
4.2 Main Results
4.2.1 Overall Performance
Here, we compare CoQ with both generic and SQL-aided table understanding methods on three datasets involving multi-hop reasoning. As shown in Tables 1 and 2, CoQ consistently achieves both higher answer accuracy and lower SQL error rates. On WikiTQ, it reaches 74.77% with GPT-3.5, a +13.66% gain over the second-best. On TabFact, it attains 92.31%, exceeding the best baseline by +10.39%. On FeTaQA, CoQ reports the highest BLEU (22.19) and ROUGE scores (R-1: 0.65, R-2: 0.42, R-L: 0.54), and the lowest error rate (7.74%). We also assess how CoQ scales with stronger models in Table 3. These results highlight the effectiveness of CoQ, which is analyzed below.
@@ 4.2.1 Overall Performance
Method LLaMA 3.1 DeepSeek-V3 GPT-4.1
Acc. Inv. Acc. Inv. Acc. Inv.
Table-to-Text 16.41 N/A 48.39 N/A 50.05 N/A
Few-Shot QA 39.94 N/A 66.07 N/A 70.97 N/A
Chain-of-Table 54.17 N/A 71.64 N/A 75.78 N/A
Basic Text-to-SQL 44.13 16.71 64.75 12.10 70.79 7.50
MAG-SQL 45.49 11.44 65.79 9.14 71.71 5.82
@@ 4.2.2 Comparison Against Generic Methods
We now provide a detailed comparison with generic table understanding methods, which rely solely on LLMs’ natural language capabilities without explicit SQL assistance. Their improvements are modest (e.g., Tree-of-Table improves accuracy by just 1.17% over Chain-of-Table on WikiTQ). This indicates a performance plateau for current methods. On the other hand, Table-to-Text methods linearize entire tables into textual descriptions to eliminate structured content. However, the generated text often exhaustively covers all table content, making it difficult for LLMs to locate key information. This leads to low accuracy and poor scalability on large tables. Empirically, the table-to-text baseline performs the worst among all methods (e.g., 16.07% on WikiTQ).
@@ 4.2.3 Comparison Against SQL Methods
We also compare CoQ with SQL-aided baselines and find that it consistently achieves higher accuracy and lower SQL error rates across all datasets and model backbones. A key limitation of existing SQL-aided methods lies in their rigid, one-shot formulation: they attempt to generate a single SQL query that directly yields the final answer. This approach works reasonably well for traditional Text-to-SQL tasks, where answers typically correspond to certain cell values. However, in table understanding, answers often require multi-step reasoning, abstraction, or comparisons across multiple rows or columns. Forcing all logic into one complex SQL query in such settings leads to bloated and error-prone programs. For example, while MAG-SQL performs well on standard Text-to-SQL benchmarks (Xie et al., 2024b), it underperforms on reasoning-intensive table understanding tasks. MAG-SQL only achieves 55.87% accuracy and a 9.48% invalid SQL rate on WikiTQ with GPT-3.5, compared to 74.77% accuracy and just 3.34% invalid SQL with CoQ. Similar trends are observed on TabFact and FeTaQA.
@@
4.3 Scalability to Stronger LLMs
@@ 4.5 Ablation Study: Deep Dive into CoQ’s Key Components and Mechanisms
We conduct an ablation study on WikiTQ using GPT-3.5, disabling one CoQ component at a time.
@@
Table 6: Comparison of the number of LLM calls per question across methods on WikiTQ dataset. CoQ maintains high performance with low LLM usage, enabled by hybrid reasoning division that avoids over-generation.
@@ 4.6 Analysis of LLM Usage
In addition to accuracy, we examine CoQ’s LLM usage by analyzing the upper bound of calls required per question (Table 6). Among all methods, CoQ offers a favorable trade-off between reasoning depth and LLM usage, requiring at most 22 calls. In contrast, other strong baselines require up to 25–100 calls due to extensive branching and fixed steps. While MAC-SQL is slightly more efficient (within 19 calls), it underperforms on complex tasks. Notably, we report upper bounds here; in practice, CoQ averages just 7.63 LLM calls per instance (Appendix E). This low average stems from our Hybrid Reasoning Division: by delegating logical reasoning to the LLM, CoQ avoids generating overly complex SQL and reduces error correction. It issues only the essential clauses needed to retrieve sufficient information, minimizing execution overhead.
@@ 6 Limitations
7 Ethics Statement
This work uses five publicly available datasets: WikiTQ, TabFact, FeTaQA, IM-TQA, and Open-WikiTable, all of which are widely used in prior research and contain no personally identifiable information. No additional data collection or human annotation was performed. Our experiments use GPT-3.5, GPT-4.1, LLaMA 2, and DeepSeek-V3 in standard inference-only settings. While we are not aware of specific ethical concerns related to our datasets or methods, we acknowledge potential risks: the reliance on pretrained LLMs may propagate social or cultural biases, and large-scale model deployment could raise environmental concerns.
@@ F.1 WikiTQ
WikiTQ (Pasupat and Liang, 2015) is a question answering dataset derived from Wikipedia, a resource with broad topical coverage. It contains 2,108 tables and 22,033 natural language questions, each paired with an answer derivable from a single table. Answers typically correspond to one or more table cells.
@@ F.2 TabFact
TabFact (Chen et al., 2020) is a table-based fact verification dataset over Wikipedia tables covering over 16 domains. Each example includes a table and a natural language statement, labeled as either entailed or refuted. It comprises 117,854 examples with binary labels.
@@ F.3 FeTaQA
FeTaQA (Nan et al., 2022) is a complex table QA dataset focused on multi-hop reasoning. Each entry includes a Wikipedia table, a natural language question, and a free-form textual answer. It focuses on free-form, multi-step reasoning beyond cell-level lookup.
@@ F.4 IM-TQA
IM-TQA (Zheng et al., 2023) is a table QA dataset that is built by collecting tables from open websites of more than 10 domains. It features complex table styles, including hierarchical, nested, and messy real-world structures. Each entry includes a structurally complex table, several natural language questions, and answers relevant to certain cells.
@@ F.5 Open-WikiTable
Open-WikiTable (Kweon et al., 2023) is a dataset for open domain question answering with complex reasoning over multi-table settings. Each entry includes a natural language question, a table related to the question, and the answer.

### v3

@@ 3.5 Balancing SQL and LLM Reasoning through Dynamic Planning
The Dynamic Planner incrementally checks each newly generated query and its execution result to decide whether the current sub-table is sufficient for LLM reasoning. If the sub-table still contains irrelevant or indirect information, the Planner selects another SQL clause and asks the SQL Generator to append it to the current query. The Planner is guided by a Sufficiency-based Early Stopping Mechanism. Together with the Answer Generator, it jointly implements the Hybrid Reasoning Division Strategy: SQL handles mechanical reasoning, while LLMs perform logical reasoning over intermediate results to produce the final answer.
@@ 3.5.1 Hybrid Reasoning Division Strategy
Existing SQL-aided approaches treat the SQL output as the final answer, requiring highly precise queries. This often leads to long, fragile SQL programs, especially when reasoning is complex or question intent is ambiguous. However, table understanding is inherently open-ended: the goal is to derive a natural language answer, not merely to retrieve an exact value. Based on these insights, we introduce a Hybrid Reasoning Division Strategy that decomposes table understanding into two stages: mechanical reasoning (e.g., filtering, arithmetic) is offloaded to SQL, whereas logical reasoning (e.g., comparison, inference) is handled by LLMs. Crucially, SQL execution results are treated as intermediate data rather than final answers, enabling LLMs to apply their stronger inference capabilities over SQL outputs to arrive at the final answer. This division leverages the respective strengths of SQL and LLMs. Our case study (Appendix A) and examples (Appendix B.2) further explain this design.
Table 1: Main results on WikiTQ and TabFact. Acc. = accuracy (%), Inv. = invalid SQL rate (%, lower is better). Underline = second-best, bold = best. Improvements are over second-best. CoQ achieves the highest accuracy with a large margin, driven by schema abstraction, clause-by-clause SQL generation, and hybrid reasoning.
@@
Generic LLM-based Table Understanding
@@
3.5.2 Sufficiency-based Early Stopping
To support this hybrid reasoning strategy, we introduce the Sufficiency-based Early Stopping Mechanism. It enables the Planner to halt clause generation once the retrieved data are sufficient. This prevents unnecessary SQL complexity and ensures that reasoning responsibilities are dynamically and appropriately balanced between SQL and LLMs. The detailed algorithm is provided in Appendix D.
3.6 Answer Generator
@@ 4 Experiments
In this section, we empirically evaluate the effectiveness of CoQ and aim to answer the following questions: RQ1. How does CoQ compare with popular table understanding methods (e.g., generic, SQL-based, and hybrid)? RQ2. How well does CoQ generalize to real-world, structurally complex tabular workloads? RQ3. How does CoQ perform relative to existing methods in terms of cost?
@@ 4.1 Experimental Setup
Datasets and Evaluation Metrics. We evaluate our CoQ framework on five widely used table understanding benchmarks: WikiTQ (pasupat2015wikitab), FeTaQA (Nan2021FeTaQAFT), TabFact (2019TabFactA), IM-TQA (zheng-etal-2023-im), and Open-WikiTable (kweon2023openwikitabledatasetopendomain). WikiTQ and FeTaQA are table QA datasets requiring short-span and free-form answers, respectively. TabFact is a fact verification task based on tables. IM-TQA features complex real-world table styles, while Open-WikiTable involves multi-table scenarios. Collectively, these datasets cover diverse reasoning types, table structures, and domains, forming a robust testbed for evaluating generalization. Dataset details are provided in Appendix F.
We use official accuracy metrics for WikiTQ, IM-TQA, and Open-WikiTable, and standard binary accuracy for TabFact. For FeTaQA, we report BLEU (papineni2002bleu) and ROUGE (lin2004rouge) scores to align with prior work. We also track the invalid SQL generation rate.
Baselines. We first compare our method against two categories of baseline approaches: (a) Generic LLM-based table understanding, including End-to-End QA, Few-Shot QA, Table-to-Text (min2024exploringimpacttabletotextmethods), Chain-of-Thought (wei2022cot), Binder (Binder), Dater (ye2023dater), Chain-of-Table (wang2024chaintab), and Tree-of-Table (ji2024treetab); (b) SQL-aided table understanding, including Basic Text-to-SQL (rajkumar2022evaluating), OpenTab (kong2024opentab), MAC-SQL (wang2025macsql), and MAG-SQL (xie2024mag). We further compare our CoQ with two more categories of baselines: (c) Hybrid table understanding, including TabSQLify (nahid2024tabsqlifyenhancingreasoningcapabilities), SynTQA (zhang2024syntqasynergistictablebasedquestion), TableMaster (cao2025tablemasterrecipeadvancetable), and H-STAR (abhyankar2025hstarllmdrivenhybridsqltext); (d) General structured data reasoning, including StructGPT (jiang2023structgptgeneralframeworklarge), QueryAgent (huang2024queryagentreliableefficientreasoning), Readi (cheng2024necessaryllmsefficientlyfaithfully), and Chain-of-Knowledge (li2024chainofknowledgegroundinglargelanguage). Results and analysis for category (d) are provided in Appendix J.2 due to page limitations.
Implementation Details. To align with baselines, we adopt GPT-3.5 and LLaMA 2 as the backbone LLMs, and additionally evaluate with the recent stronger models LLaMA 3.1, DeepSeek-V3, and GPT-4.1. Model configurations are detailed in Appendix H. Prompts include few-shot examples sampled from the training set, with illustrative cases shown in Appendix K.
@@
4.2 Main Results
4.2.1 Overall Performance
Here, we compare CoQ with both generic and SQL-aided table understanding methods on three datasets involving multi-hop reasoning. As shown in Tables 1 and 2, CoQ consistently achieves both higher answer accuracy and lower SQL error rates. On WikiTQ, it reaches 74.77% with GPT-3.5, a +13.66% gain over the second-best. On TabFact, it attains 92.31%, exceeding the best baseline by +10.39%. On FeTaQA, CoQ reports the highest BLEU (22.19) and ROUGE scores (R-1: 0.65, R-2: 0.42, R-L: 0.54), and the lowest error rate (7.74%). We also assess how CoQ scales with stronger models in Appendix J.1 and further analyze a challenging subset of WikiTQ in Appendix J.3. These results highlight the effectiveness of CoQ, which is analyzed below.
Table 3: Results of CoQ and hybrid baselines on WikiTQ and TabFact datasets (GPT-3.5). CoQ achieves the highest accuracy, attributed to its schema abstraction and fine-grained SQL generation control.
Method WikiTQ TabFact
TabSQLify 64.7 79.5
TableMaster 68.2 83.7
H-STAR 69.6 85.0
SynTQA 70.4 N/A
Chain-of-Query (Ours) 74.8 [-1pt] (+4.4) 92.3 [-1pt] (+7.3)
@@ 4.2.2 Comparison Against Generic Methods
We now provide a detailed comparison with generic LLM-based table understanding methods, which rely solely on LLMs’ natural language capabilities without explicit SQL assistance. Their improvements are modest (e.g., Tree-of-Table improves accuracy by just 1.17% over Chain-of-Table on WikiTQ). This indicates a performance plateau for current methods. On the other hand, Table-to-Text methods linearize entire tables into textual descriptions to eliminate structured content. However, the generated text often exhaustively covers all table content, making it difficult for LLMs to locate key information. This leads to low accuracy and poor scalability on large tables. Empirically, the table-to-text baseline performs the worst among all methods (e.g., 16.07% on WikiTQ).
@@ 4.2.3 Comparison Against SQL Methods
We also compare CoQ with SQL-aided baselines and find that it consistently achieves higher accuracy and lower SQL error rates across all datasets and model backbones. A key limitation of existing SQL-aided methods lies in their rigid, one-shot formulation: they attempt to generate a single SQL query that directly yields the final answer. This approach works reasonably well for traditional Text-to-SQL tasks, where answers typically correspond to certain cell values. However, in table understanding, answers often require multi-step reasoning, abstraction, or comparisons across multiple rows or columns. Forcing all logic into one complex SQL query in such settings leads to bloated and error-prone programs. For example, while MAG-SQL performs well on standard Text-to-SQL benchmarks (xie2024mag), it underperforms on reasoning-intensive table understanding tasks. MAG-SQL only achieves 55.87% accuracy and a 9.48% invalid SQL rate on WikiTQ with GPT-3.5, compared to 74.77% accuracy and just 3.34% invalid SQL with CoQ. Similar trends are observed on TabFact and FeTaQA.
@@
4.3 Comparison Against Hybrid Methods
Here, we compare the performance of CoQ with hybrid methods on the WikiTQ and TabFact datasets using GPT-3.5. As shown in Table 3, CoQ consistently achieves the best performance across both datasets.
Notably, although our work shares the same hybrid philosophy, it differs significantly in granularity and robustness. Unlike these approaches, which rely on simple SQL operations and one-shot SQL generation without invalid-query correction, CoQ employs a clause-by-clause validated SQL generation process with sufficiency checks and rollback mechanisms. This ensures that the generated SQL is both valid and adaptively complex for deeper mechanical reasoning. Logical inference is then handled by the LLM over the intermediate SQL outputs. This design achieves a more fine-grained separation between mechanical and logical reasoning while ensuring robustness. The key to this robust, fine-grained separation lies in dynamic planning, where the Planner adaptively adjusts the SQL generation depth depending on the input complexity.
@@ 4.5 Ablation Study: Deep Dive into CoQ’s Key Components and Mechanisms
We conduct an ablation study on WikiTQ using GPT-3.5, disabling one CoQ component at a time. A more detailed analysis is provided in Appendix J.4.
@@
Table 6: Comparison of the theoretical number of LLM calls per question across methods on WikiTQ dataset.
@@ 4.6 Analysis of LLM Usage
In addition to accuracy, we examine CoQ’s LLM usage by analyzing the theoretical upper bound of calls required per question (Table 6). Among all methods, CoQ offers a favorable trade-off between reasoning depth and LLM usage, requiring at most 22 calls. In contrast, other strong baselines require up to 25–100 calls due to extensive branching and fixed steps. While MAC-SQL is slightly more efficient (within 19 calls), it underperforms on complex tasks. Notably, we report upper bounds here; in practice, CoQ averages just 7.63 LLM calls per instance (Appendix E). This low average stems from our Hybrid Reasoning Division: by delegating logical reasoning to the LLM, CoQ avoids generating overly complex SQL and reduces error correction. It issues only the essential clauses needed to retrieve sufficient information, minimizing execution overhead. A detailed step-wise breakdown of LLM calls and an analysis of token-level cost are provided in Appendix J.5.
@@ 5 Conclusion
Ethics Statement
This work uses five publicly available datasets: WikiTQ, TabFact, FeTaQA, IM-TQA, and Open-WikiTable, all of which are widely used in prior research and contain no personally identifiable information. No additional data collection or human annotation was performed. Our experiments use GPT-3.5, GPT-4.1, LLaMA 2, LLaMA 3.1, and DeepSeek-V3 in standard inference-only settings. While we are not aware of specific ethical concerns related to our datasets or methods, we acknowledge potential risks: the reliance on pretrained LLMs may propagate social or cultural biases, and large-scale model deployment could raise environmental concerns.
Acknowledgments
We thank all of our team members for their dedicated efforts and valuable discussions throughout this project, especially during the author–reviewer discussion phase. We thank the authors of Chain-of-Table, MAG-SQL, and OpenTab for sharing their code. Lastly, we sincerely appreciate all reviewers for their invaluable feedback, insightful suggestions, and positive remarks about our work.
@@ A.1 Hybrid Reasoning
A.2 Handling String Variations
For noisy entries such as "New York" vs. "new york*" or "Dallas Cowboys" vs. "Dallas Cowboys*", our framework addresses the issue through fuzzy matching in SQL generation and refinement.
For example, in questions such as "How many games did the cowboys play?", the LLM generates SQL queries like SELECT ... WHERE team LIKE "%Dallas Cowboys%".
Appendix K.11 shows our prompt for correcting invalid WHERE clauses, and its "Constraints" section encourages the LLM to relax filtering conditions, including using fuzzy matching "IN" and "LIKE".
This mechanism ensures that variations with minor inconsistencies are still matched. While fuzzy matching cannot cover all possible noise patterns, in practice it significantly reduces such errors and allows CoQ to remain robust to imperfect table entries.
A.3 Handling Special Rows
The key challenge of handling such rows lies in their special values. In our framework, these rows are addressed through the integration of the Natural-Language-Style Table Schema and Hybrid Reasoning.
The schema provides representative value examples for each column, which are typically normal values randomly sampled from the table. Our SQL-based mechanical reasoning does not directly yield the final answer but instead produces an intermediate sub-table. This allows the LLM to compare the schema examples with the SQL results and detect anomalies such as special values from the aggregate rows.
For instance, Figure 4 shows a table with special rows studied by liu2023rethinkingtabulardataunderstanding.In their framework, the Python-based method attempts to directly return the exact number, counting the aggregate row by mistake. However, our method provides the LLM with intermediate information as illustrated in Figure 5.
Figure 4: An example of a table containing a special row, where the last row includes "Total".
Figure 5: Illustration of CoQ’s intermediate information in this case.
In this example, the LLM can distinguish the normal entry "2004" from the special entry "total" and correctly infer that the result contains an aggregate row, thus avoiding misinterpretation.
@@ Appendix D Algorithm of Sufficiency-based Early Stopping Mechanism
We present the pseudocode for the Sufficiency-Based Early Stopping Mechanism, which allows the Dynamic Planner to terminate clause generation once sufficient information has been retrieved.
@@ F.1 WikiTQ
WikiTQ (pasupat2015wikitab) is a question answering dataset derived from Wikipedia, a resource with broad topical coverage. It contains 2,108 tables and 22,033 natural language questions, each paired with an answer derivable from a single table. Answers typically correspond to one or more table cells.
@@ F.2 TabFact
TabFact (2019TabFactA) is a table-based fact verification dataset over Wikipedia tables covering over 16 domains. Each example includes a table and a natural language statement, labeled as either entailed or refuted. It comprises 117,854 examples with binary labels.
@@ F.3 FeTaQA
FeTaQA (Nan2021FeTaQAFT) is a complex table QA dataset focused on multi-hop reasoning. Each entry includes a Wikipedia table, a natural language question, and a free-form textual answer. It focuses on free-form, multi-step reasoning beyond cell-level lookup.
@@ F.4 IM-TQA
IM-TQA (zheng-etal-2023-im) is a table QA dataset that is built by collecting tables from open websites of more than 10 domains. It features complex table styles, including hierarchical, nested, and messy real-world structures. Each entry includes a structurally complex table, several natural language questions, and answers relevant to certain cells.
@@ F.5 Open-WikiTable
Open-WikiTable (kweon2023openwikitabledatasetopendomain) is a dataset for open domain question answering with complex reasoning over multi-table settings. Each entry includes a natural language question, a table related to the question, and the answer.